# A Test of Substellar Evolutionary Models with High-Precision Ages from Asteroseismology and Gyrochronology for the Benchmark System HR 7672AB

### v1

@@ 2.2 Dynamical Orbit with NIRC2 and Radial Velocities
2001.64 0.786 0.006 157.9 0.5 Liu2002
2001.94 0.794 0.005 157.3 0.6 Liu2002
2002.54 0.788 0.006 156.6 0.9 Boccaletti2003
2006.69 0.750 0.080 155.0 5.0 Serabyn2009
2007.73 0.742 0.035 151.8 2.9 Crepp2012
2011.37 0.519 0.006 147.1 0.5 Crepp2012
2024.78 0.852 0.007 329.5 1.1 §2.2
@@
Our new measurement, along with all available measurements, are summarized in Table 3. In addition, HR 7672 A has absolute astrometry from both Hipparcos (ESA1997) and Gaia (GaiaCollaboration2022), allowing us to measure its proper motion acceleration over a 20-year baseline. These data are listed in Table 4. The California Planet Search (CPS) has acquired RVs of HR 7672A from 1970 to 2024 using Lick/APF and Keck/HIRES instruments (Fischer2014; Rosenthal2021). All RVs are compiled in Table 5.
We modeled the dynamical orbit by jointly fitting the Hipparcos and Gaia astrometry, direct imaging measurements, and RV data. We used open source package orvara (Brandt2021b), which performs a parallel-tempering Markov Chain Monte Carlo (MCMC) fitting. In total, our analysis used 16 free parameters. Two of them are the masses of the primary star ($M_{\rm{A}}$) and brown dwarf companion ($M_{\rm{B}}$). Six orbital parameters define the orbit of companion: semi-major axis ($a_{B}$), inclination ($i_{B}$), longitude of the ascending node ($\Omega_{B}$), mean longitude at a reference epoch ($t_{\rm ref}$) of 2455197.5 JD ($\lambda_{\rm ref}$), and the eccentricity ($e_{B}$) and the argument of periastron ($\omega_{B}$) in the form of $e_{B}\sin\omega_{B}$ and $e_{B}\cos\omega_{B}$. We also included four parameters to fit the zero-point for RV data from different instruments. Finally, we have four parameters for the intrinsic jitter of RV data from different instruments. Because HIRES underwent an instrument upgrade in August 2014, we adopted separate radial velocity zero points and jitters for data taken before and after the upgrade. The likelihood was computed by comparing the observed separations, position angles, absolute astrometry, and radial velocities with those predicted from a synthetic orbit, under the assumption of Gaussian measurement errors.
We used 100 walkers to sample our model with a total of $2.5\times 10^{5}$ steps. The chain converged within the first 30%, which we discarded as the burn-in portion. The best-fitting results to the astrometric and RV data are presented in §A. In particular, we determined a primary-star mass of $1.11\pm 0.02$ $\text{M}_{\odot}$, and a companion mass of $75.4\pm 0.7$ $\text{M}_{\text{Jup}}$. For comparison, Crepp2012 reported a companion mass of $68.7\pm 2.8$$\text{M}_{\text{Jup}}$, Brandt2019 obtained $72.7\pm 0.8$$\text{M}_{\text{Jup}}$, Feng2021 found $72.8\pm 6.1$$\text{M}_{\text{Jup}}$, and An2025 derived $71.43\pm 1.05$$\text{M}_{\text{Jup}}$ (astrometry only). The difference primarily results from the inclusion of updated radial velocity and astrometric data.
2.3 Rotation and Activity
@@
To obtain a rotation period from the S index, which is modulated by the star’s rotation, we modeled the time series using a hierarchical model and Gaussian processes (celerite2). Each time series was assigned a single rotation period, with all periods assumed to follow a normal distribution with a common mean and a standard deviation, using the rotation kernel term in celerite2. This rotation kernel is a sum of two simple harmonic oscillator kernels (c.f., celerite2). A hierarchical model is well suited in this situation because of varying systematics, noise levels, and signal amplitudes across instruments. Using this model, we determined a rotation period of ${P_{\rm rot}}=14.8\pm 1.3$ d. This value agrees well with the rotation periods derived in previous work such as 13.94 d from Donahue1996 from Mount Wilson data alone, and 13.95 d from Messina2001 from V-band photometry, and 14.54 d from Keck/HIRES.
@@
The average S-index of 0.19 corresponds to a $\log R^{\prime}_{\rm HK}$ value of 4.74-4.74 (BoroSaikia2018), placing the star firmly within an active regime where standard magnetic spin-down remains applicable. Weakened magnetic braking is only expected to commence when $\log R^{\prime}_{\rm HK}$ reaches 5.0-5.0, or equivalently, at a Rossby number of $0.9{{\rm Ro}}_{\odot}$ (vanSaders2016-wmb; saunders2024-ro).
Rotation lifts the degeneracy between modes with different $m$ but the same $n$ and $\ell$. In this case, the frequencies are given by $\nu_{n\ell m}\approx\nu_{n\ell}+m\nu_{\rm rot}$, where $\nu_{\rm rot}$ is the rotational frequency (Ledoux1951). The maximum possible splitting occurs between the $m=$-$\ell$ and $m=\ell$ components and is thus $2\ell\nu_{\rm rot}$, which is reached when the rotation axis is viewed edge-on (Gizon2003). Given the measured rotation period, the splittings are 1.56, 3.13, and 4.69 $\mu{\rm Hz}$ for $\ell=1,2,3$ modes, respectively. Indeed, some broadening is observed in the ridges formed by non-radial modes in Figure 2(b), which could plausibly be attributed to this rotational splitting.
2.4 Effective Temperature and Metallicity
To determine the spectroscopic parameters, including effective temperature $T_{\text{eff}}$ and metallicity $\rm[M/H]$}, we adopted the values from Rosenthal2021, derived from Keck/HIRES spectra. They are 5932 K and 0.051 dex, respectively. We also compiled $T_{\text{eff}}$ and $\rm[M/H]$} values in literature using the SIMBAD astronomical database. These values exhibit a spread of around 100 K in $T_{\text{eff}}$ and 0.1 dex in $\rm[M/H]$}, which we adopt as the representative uncertainties. These are consistent with typical spectroscopic uncertainties for solar-type stars (Serenelli2017; Tayar2022).
2.5 Luminosity
We used isoclassify (Huber2017; Berger2020) to calculate the bolometric flux based on apparent magnitudes from the Tycho-2, Gaia, and 2MASS photometric systems. Combined with the Gaia DR3 distance (bj2021), this yields an estimate of the stellar luminosity. Depending on the band used, the derived luminosity ranges from 1.0 to 1.75 $\text{L}_{\odot}$, with typical formal uncertainties of approximately 0.04 $\text{L}_{\odot}$. The large spread reflects systematic issues such as saturation in bright-star photometry, uncertainties in bolometric corrections, and photometric zero points (Riello2021). The Gaia BP and RP bands are the most discrepant. Among the available bands, the luminosity value based on the 2MASS $K_{s}$ band appears most consistent with that derived from asteroseismic modeling in §3.1. However, we did not use it as a modeling constraint due to the inconsistency across photometric bands. All stellar parameters are summarized in Table 7.
@@ 2.5 Luminosity
Pulsation code GYRE v7.1 GYRE v5 ADIPLS v0.4 antia1994 GYRE v7.2 ADIPLS
@@
Atmospheric boundary condition hauschildt1999a; hauschildt1999b castelli2003 ks1966 eddington1926 eddington1926 trampedach2014 eddington1926
Metal mixture agss09 agss09 agss09 gs98 gs98 gs98
Nuclear reaction rate jinareaclib2010 nacre1999 jinareaclib2010 nacre1999 nacre1999 formicola2004 hammer2005 adelberger1998 formicola2004 nacre1999 formicola2004 hammer2005 nacre1999
Mixing length formulation henyey1965 cox1968 bv1958 bv1958 cox1968 bv1958
@@
Convective overshoot herwig2000 $f_{\rm ov,shell}=0.0174$ None None None None None
Extra-mixing None Gravitational settling Turbulent mixing thoul1994 thoul1994 thoul1994 Diffusive mixing Michaud1993
Surface correction bg14 two-term +liyg2023 ensemble correction sonoi2015 bg14 cubic-term bg14 two-term bg14 two-term + Roxburgh2016 $\epsilon$-matching Kjeldsen2008
@@
$K_{s}$-band magnitude 4.388 $\pm$ 0.027 Cutri2003
@@
Mass ($\text{M}_{\odot}$) 1.111 $\pm$ 0.017 1,2 §2.2
Mass ($\text{M}_{\text{Jup}}$; HR 7672 B) 75.39 $\pm$ 0.67 §2.2
@@
Distance (pc) 11.758 $\pm$ 0.014 bj2021
Luminosity $L_{\star}$ ($\text{L}_{\odot}$) 1.224 $\pm$ 0.043 §2.5
@@
Effective temperature $T_{\text{eff}}$ (K) 5932 $\pm$ 81 1,2 Rosenthal2021
Metallicity $\rm[Fe/H]$} (dex) 0.051 $\pm$ 0.057 1,2 Rosenthal2021
@@
$\nu_{n,\ell}$ ($\mu{\rm Hz}$) Table 2 1 §2.1
$\epsilon$ $1.335\pm 0.023$ §2.1
$\Delta\nu$ ($\mu{\rm Hz}$) $132.806\pm 0.126$ §2.1
$\delta\nu_{0,1}$ ($\mu{\rm Hz}$) $5.47\pm 0.69$ §2.1
$\delta\nu_{0,2}$ ($\mu{\rm Hz}$) $12.53\pm 0.67$ §2.1
$\delta\nu_{0,3}$ ($\mu{\rm Hz}$) $22.87\pm 1.17$ §2.1
$\nu_{\text{max}}$ ($\mu{\rm Hz}$) $3245\pm 75$ §2.1
$A_{\text{osc}}$ (cm s-1) $16.9\pm 0.6$ §5
$A_{\text{osc}}$ (ppm) $3.4\pm 0.9$ §5
Mass ($M_{\odot}$) $1.114$ $\pm$ $0.013$ (stat) $\pm$ $0.002$ (sys) §3.1
Radius ($R_{\odot}$) $1.051$ $\pm$ $0.004$ (stat) $\pm$ $0.010$ (sys) §3.1
Luminosity ($L_{\odot}$) $1.25$ $\pm$ $0.05$ (stat) $\pm$ $0.01$ (sys) §3.1
Asteroseismic age (Gyr) $1.87$ $\pm$ $0.56$ (stat) $\pm$ $0.32$ (sys) §3.1
Density ($\rho_{\odot}$) $0.960$ $\pm$ $0.002$ (stat) $\pm$ $0.026$ (sys) §3.1
$\log g$ (cgs; dex) $4.442$ $\pm$ $0.002$ (stat) $\pm$ $0.010$ (sys) §3.1
@@
Rotation period $P_{\rm rot}$ (d) 14.8 $\pm$ 1.3 2 §2.3
Rotational age (Gyr) 2.58 $\pm$ 0.47 §3.2
@@
Asteroseismic + rotational joint age (Gyr) 2.26 $\pm$ 0.40 §3.3
@@ 3.1 Asteroseismic Modeling
Stellar evolution codes used by the six teams included MESA (mesa2011; mesa2013; mesa2015; mesa2018; mesa2019; mesa2023; Moedas2024), GARSTEC (garstec2008), YREC (yrec2008), and ASTEC (ChristensenDalsgaard2008a). Pulsation codes used for calculating oscillation frequencies included GYRE (gyre2013), ADIPLS (adipls2008), and the code used by antia1994.
The input physics used by the six teams also differed. Treatments of atmospheric boundary conditions included the Eddington (eddington1926) and Krishna-Swamy (ks1966) $T$-$\tau$ relations, and pre-computed photosphere tables (hauschildt1999a; hauschildt1999b; castelli2003; trampedach2014). Choices of nuclear reaction rates varied from jinareaclib2010, nacre1999, formicola2004, hammer2005, and adelberger1998. The mixing length formulations included henyey1965, cox1968 and bv1958, with some teams using solar-calibrated mixing length parameters, others treating it as a free variable, and still others using calibrated prescriptions for varying it. Two main metal mixtures were used: those of agss09 and of gs98. Mass, metallicity, and initial helium abundance were all treated as free parameters by all teams.
Different teams employed different correction procedures for near-surface modeling errors. Several of the teams used empirical correction formulae (Kjeldsen2008; bg14), where an offset is applied to the mode frequencies to account for the effects of the surface term, before the corrected mode frequencies are used to construct a conventional likelihood function. These empirical formulae admit additional free parameters, which may either be fitted against the data, or potentially be externally calibrated against 3D models (sonoi2015) or from an empirical ensemble (liyg2023). Other teams employed the nonparametric $\epsilon$-matching technique (Roxburgh2016), where combinations of mode frequencies are computed from both the model and observed set, chosen so that the likelihood function evaluated from these combinations is insensitive to the near-surface structure of the star (Ong2021-surferr). Table 6 lists the detailed configurations as well as the derived stellar parameters by each team.
@@
3.2 Rotational Modeling
To model the star based on its rotation period using gyrochronology, we require rotation period predictions from stellar models. For this purpose, we used Team 1’s stellar models in combination with the rotevol code (vanSaders2013-sg; somers2017). Angular momentum loss was modeled using a calibrated prescription, based on open cluster data and asteroseismic field stars, following the exact configuration described in Liyg2025.
We briefly summarize the calibration procedure used in the angular momentum loss modeling. Stellar models were initialized with a rotation period of $P_{\rm disk}=4$ days and a disk-locking timescale of $\tau_{\rm disk}=10$ Myr, following the setup of chiti2024-wd. We note that these initial conditions does not influence the rotation–age relation once convergence has been achieved. Angular momentum loss was modeled using the prescription from vanSaders2016-wmb, which relates the loss rate to the stellar rotation rate and convective overturn timescale, scaled by a braking efficiency parameter $f_{K}$. When the Rossby number exceeds a critical threshold ${{\rm Ro}}_{\rm crit}$, the star enters a weakened magnetic braking regime, and angular momentum loss is no longer applied. This is less relevant in our case because the star has not entered into that phase. Stellar rotation was modeled using a two-zone framework in which the radiative core and convective envelope rotate as solid bodies but exchange angular momentum over a coupling timescale $\tau_{\rm ce}$. This timescale is assumed to scale with stellar mass as $\tau_{\rm ce}/\tau_{\rm ce,\odot}=(M/M_{\odot})^{-\alpha_{\rm ce}}$, where $\tau_{\rm ce,\odot}\approx 22$ Myr (sl20). The free parameters in the model — $f_{K}$, ${{\rm Ro}}_{\rm crit}$, and $\alpha_{\rm ce}$ — were calibrated to reproduce the rotation periods of stars in a set of well-characterized open clusters ranging from the 120 Myr Pleiades to the 4 Gyr M67 (rebull2016-pleiades; douglas2017-praesepe; douglas2019-praesepe; curtis2019-6811; meibom2015-6819; curtis2020-147; barnes2016-m67; dungee2022-m67; long2023-cl), asteroseismic field stars (hall2021; silvaaguirre2015-kages; silvaaguirre2017-legacy), and the Sun (Bahcall1995) at their respective ages. The $T_{\text{eff}}$ and rotation period of HR 7672A place it in a regime where gyrochronology is well calibrated.

### v2

@@ II.3 Dynamical Orbit with NIRC2 and Radial Velocities
2001.64 0.786 0.006 157.9 0.5 M. C. Liu et al. (2002a)
2001.94 0.794 0.005 157.3 0.6 M. C. Liu et al. (2002a)
2002.54 0.788 0.006 156.6 0.9 A. Boccaletti et al. (2003)
2006.69 0.750 0.080 155.0 5.0 E. Serabyn et al. (2009)
2007.73 0.742 0.035 151.8 2.9 J. R. Crepp et al. (2012)
2011.37 0.519 0.006 147.1 0.5 J. R. Crepp et al. (2012)
2024.78 0.852 0.007 329.5 1.1 §II.3
@@
Our new measurement, along with all available measurements, are summarized in Table 3. In addition, HR 7672 A has absolute astrometry from both Hipparcos (ESA, 1997) and Gaia (Gaia Collaboration, 2022), allowing us to measure its proper motion acceleration over a 20-year baseline. These data are listed in Table 4. The California Planet Search (CPS) has acquired RVs of HR 7672A from 1970 to 2024 using Lick/APF and Keck/HIRES instruments (D. A. Fischer et al., 2014; L. J. Rosenthal et al., 2021). All RVs are compiled in Table 5.
We modeled the dynamical orbit by jointly fitting the Hipparcos and Gaia astrometry, direct imaging measurements, and RV data. We used open source package orvara (T. D. Brandt et al., 2021), which performs a parallel-tempering Markov Chain Monte Carlo (MCMC) fitting. In total, our analysis used 16 free parameters. Two of them are the masses of the primary star ($M_{\rm{A}}$) and brown dwarf companion ($M_{\rm{B}}$). We adopted a log-uniform prior on $M_{\rm{A}}$. Six orbital parameters define the orbit of companion: semi-major axis ($a_{B}$), inclination ($i_{B}$), longitude of the ascending node ($\Omega_{B}$), mean longitude at a reference epoch ($t_{\rm ref}$) of 2455197.5 JD ($\lambda_{\rm ref}$), and the eccentricity ($e_{B}$) and the argument of periastron ($\omega_{B}$) in the form of $e_{B}\sin\omega_{B}$ and $e_{B}\cos\omega_{B}$. We also included four parameters to fit the zero-point for RV data from different instruments. Finally, we have four parameters for the intrinsic jitter of RV data from different instruments. Because HIRES underwent an instrument upgrade in August 2014, we adopted separate radial velocity zero points and jitters for data taken before and after the upgrade. The likelihood was computed by comparing the observed separations, position angles, absolute astrometry, and radial velocities with those predicted from a synthetic orbit, under the assumption of Gaussian measurement errors.
We used 100 walkers to sample our model with a total of $2.5\times 10^{5}$ steps. The chain converged within the first 30%, which we discarded as the burn-in portion. The best-fitting results to the astrometric and RV data are presented in §A. In particular, we determined a primary-star mass of $1.11\pm 0.02$ $\text{M}_{\odot}$, and a companion mass of $75.4\pm 0.7$ $\text{M}_{\text{Jup}}$. For comparison, J. R. Crepp et al. (2012) reported a companion mass of $68.7\pm 2.8$$\text{M}_{\text{Jup}}$, T. D. Brandt et al. (2019a) obtained $72.7\pm 0.8$$\text{M}_{\text{Jup}}$, F. Feng et al. (2021) found $72.8\pm 6.1$$\text{M}_{\text{Jup}}$, and Q. An et al. (2025a) derived $71.43\pm 1.05$$\text{M}_{\text{Jup}}$. The difference primarily reflects the inclusion of updated radial-velocity and astrometric data. Q. An et al. (2025a) also adopted an informative prior on the host star (a normal prior with a small standard deviation), which can significantly shift the inferred companion mass. Because both component masses are independently constrained in this system, we recommend using uninformative priors, as in this work, T. D. Brandt et al. (2019a), and F. Feng et al. (2021).
II.4 Rotation and Activity
@@
To obtain a rotation period from the S index, which is modulated by the star’s rotation, we modeled the time series using a hierarchical model and Gaussian processes (D. Foreman-Mackey, 2018). Each time series was assigned a single rotation period, with all periods assumed to follow a normal distribution with a common mean and a standard deviation, using the rotation kernel term in celerite2. This rotation kernel is a sum of two simple harmonic oscillator kernels (c.f., D. Foreman-Mackey 2018). A hierarchical model is well suited in this situation because of varying systematics, noise levels, and signal amplitudes across instruments. Using this model, we determined a rotation period of ${P_{\rm rot}}=14.8\pm 1.3$ d. This value agrees well with the rotation periods derived in previous work such as 13.94 d from R. A. Donahue et al. (1996) from Mount Wilson data alone, and 13.95 d from S. Messina et al. (2001) from V-band photometry, and 14.54 d from Keck/HIRES.
@@
The average S-index of 0.19 corresponds to a $\log R^{\prime}_{\rm HK}$ value of 4.74-4.74 (S. Boro Saikia et al., 2018), placing the star firmly within an active regime where standard magnetic spin-down remains applicable. Weakened magnetic braking is only expected to commence when $\log R^{\prime}_{\rm HK}$ reaches 5.0-5.0, or equivalently, at a Rossby number of $0.9{{\rm Ro}}_{\odot}$ (J. L. van Saders et al., 2016; N. Saunders et al., 2024).
Rotation lifts the degeneracy between modes with different $m$ but the same $n$ and $\ell$. In this case, the frequencies are given by $\nu_{n\ell m}\approx\nu_{n\ell}+m\nu_{\rm rot}$, where $\nu_{\rm rot}$ is the rotational frequency (P. Ledoux, 1951). The maximum possible splitting occurs between the $m=$-$\ell$ and $m=\ell$ components and is thus $2\ell\nu_{\rm rot}$, which is reached when the rotation axis is viewed edge-on (L. Gizon & S. K. Solanki, 2003). Given the measured rotation period, the splittings are 1.56, 3.13, and 4.69 $\mu{\rm Hz}$ for $\ell=1,2,3$ modes, respectively. Indeed, some broadening is observed in the ridges formed by non-radial modes in Figure 2(b), which could plausibly be attributed to this rotational splitting.
II.5 Effective Temperature and Metallicity
To determine the spectroscopic parameters, including effective temperature $T_{\text{eff}}$ and metallicity $\rm[M/H]$}, we adopted the values from L. J. Rosenthal et al. (2021), derived from Keck/HIRES spectra. They are 5932 K and 0.051 dex, respectively. We also compiled $T_{\text{eff}}$ and $\rm[M/H]$} values in literature using the SIMBAD astronomical database. These values exhibit a spread of around 100 K in $T_{\text{eff}}$ and 0.1 dex in $\rm[M/H]$}, which we adopt as the representative uncertainties. These are consistent with typical spectroscopic uncertainties for solar-type stars (A. Serenelli et al., 2017; J. Tayar et al., 2022).
II.6 Luminosity
We used isoclassify (D. Huber et al., 2017; T. A. Berger et al., 2020) to calculate the bolometric flux based on apparent magnitudes from the Tycho-2, Gaia, and 2MASS photometric systems. Combined with the Gaia DR3 distance (C. A. L. Bailer-Jones et al., 2021), this yields an estimate of the stellar luminosity. Depending on the band used, the derived luminosity ranges from 1.0 to 1.75 $\text{L}_{\odot}$, with typical formal uncertainties of approximately 0.04 $\text{L}_{\odot}$. The large spread reflects systematic issues such as saturation in bright-star photometry, uncertainties in bolometric corrections, and photometric zero points. The Gaia BP and RP bands are the most discrepant, likely because the systematic errors for bright stars are not yet well calibrated for those Gaia bands (M. Riello et al., 2021). Among the available bands, the luminosity value based on the 2MASS $K_{s}$ band appears most consistent with that derived from asteroseismic modeling in §III.1. However, we did not use it as a modeling constraint due to the inconsistency across photometric bands. All stellar parameters are summarized in Table 7.
@@ II.6 Luminosity
Pulsation code GYRE v7.1 GYRE v5 ADIPLS v0.4 H. M. Antia & S. Basu (1994) GYRE v7.2 ADIPLS
@@
Atmospheric boundary condition P. H. Hauschildt et al. (1999a, b) F. Castelli & R. L. Kurucz (2003) K. S. Krishna Swamy (1966) A. S. Eddington (1926) A. S. Eddington (1926) R. Trampedach et al. (2014) A. S. Eddington (1926)
Metal mixture M. Asplund et al. (2009) M. Asplund et al. (2009) M. Asplund et al. (2009) N. Grevesse & A. J. Sauval (1998) N. Grevesse & A. J. Sauval (1998) N. Grevesse & A. J. Sauval (1998)
Nuclear reaction rate R. H. Cyburt et al. (2010) C. Angulo et al. (1999) R. H. Cyburt et al. (2010) C. Angulo et al. (1999) C. Angulo et al. (1999) A. Formicola et al. (2004) J. W. Hammer et al. (2005) E. G. Adelberger et al. (1998) A. Formicola et al. (2004) C. Angulo et al. (1999) A. Formicola et al. (2004) J. W. Hammer et al. (2005) C. Angulo et al. (1999)
Mixing length formulation L. Henyey et al. (1965) J. P. Cox & R. T. Giuli (1968) E. Böhm-Vitense (1958) E. Böhm-Vitense (1958) J. P. Cox & R. T. Giuli (1968) E. Böhm-Vitense (1958)
@@
Convective overshoot F. Herwig (2000) $f_{\rm ov,shell}=0.0174$ None None None None None
Extra-mixing None Gravitational settling Turbulent mixing A. A. Thoul et al. (1994) A. A. Thoul et al. (1994) A. A. Thoul et al. (1994) Diffusive mixing G. Michaud & C. R. Proffitt (1993)
Surface correction W. H. Ball & L. Gizon (2014) two-term +Y. Li et al. (2023a) ensemble correction T. Sonoi et al. (2015) W. H. Ball & L. Gizon (2014) cubic-term W. H. Ball & L. Gizon (2014) two-term W. H. Ball & L. Gizon (2014) two-term + I. W. Roxburgh (2016) $\epsilon$-matching H. Kjeldsen et al. (2008)
@@
$K_{s}$-band magnitude 4.388 $\pm$ 0.027 R. M. Cutri et al. (2003)
@@
Mass ($\text{M}_{\odot}$) 1.111 $\pm$ 0.017 1,2 §II.3
Mass ($\text{M}_{\text{Jup}}$; HR 7672 B) 75.39 $\pm$ 0.67 §II.3
@@
Distance (pc) 11.758 $\pm$ 0.014 C. A. L. Bailer-Jones et al. (2021)
Luminosity $L_{\star}$ ($\text{L}_{\odot}$) 1.224 $\pm$ 0.043 §II.6
@@
Effective temperature $T_{\text{eff}}$ (K) 5932 $\pm$ 81 1,2 L. J. Rosenthal et al. (2021)
Metallicity $\rm[Fe/H]$} (dex) 0.051 $\pm$ 0.057 1,2 L. J. Rosenthal et al. (2021)
@@
$\nu_{n,\ell}$ ($\mu{\rm Hz}$) Table 2 1 §II.1
$\epsilon$ $1.335\pm 0.023$ §II.1
$\Delta\nu$ ($\mu{\rm Hz}$) $132.806\pm 0.126$ §II.1
$\delta\nu_{0,1}$ ($\mu{\rm Hz}$) $5.47\pm 0.69$ §II.1
$\delta\nu_{0,2}$ ($\mu{\rm Hz}$) $12.53\pm 0.67$ §II.1
$\delta\nu_{0,3}$ ($\mu{\rm Hz}$) $22.87\pm 1.17$ §II.1
$\nu_{\text{max}}$ ($\mu{\rm Hz}$) $3245\pm 75$ §II.1
$A_{\text{osc}}$ (cm s-1) $16.9\pm 0.6$ §II.2
$A_{\text{osc}}$ (ppm) $3.4\pm 0.9$ §II.2
Mass ($M_{\odot}$) $1.114$ $\pm$ $0.013$ (stat) $\pm$ $0.002$ (sys) §III.1
Radius ($R_{\odot}$) $1.051$ $\pm$ $0.004$ (stat) $\pm$ $0.010$ (sys) §III.1
Luminosity ($L_{\odot}$) $1.25$ $\pm$ $0.05$ (stat) $\pm$ $0.01$ (sys) §III.1
Asteroseismic age (Gyr) $1.87$ $\pm$ $0.56$ (stat) $\pm$ $0.32$ (sys) §III.1
Density ($\rho_{\odot}$) $0.960$ $\pm$ $0.002$ (stat) $\pm$ $0.026$ (sys) §III.1
$\log g$ (cgs; dex) $4.442$ $\pm$ $0.002$ (stat) $\pm$ $0.010$ (sys) §III.1
@@
Rotation period $P_{\rm rot}$ (d) 14.8 $\pm$ 1.3 2 §II.4
Rotational age (Gyr) 2.58 $\pm$ 0.47 §III.2
@@
Asteroseismic + rotational joint age (Gyr) 2.26 $\pm$ 0.40 §III.3
@@ III.1 Asteroseismic Modeling
Stellar evolution codes used by the six teams included MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023; N. Moedas et al., 2024), GARSTEC (A. Weiss & H. Schlattl, 2008), YREC (P. Demarque et al., 2008), and ASTEC (J. Christensen-Dalsgaard, 2008a). Pulsation codes used for calculating oscillation frequencies included GYRE (R. H. D. Townsend & S. A. Teitler, 2013), ADIPLS (J. Christensen-Dalsgaard, 2008b), and the code used by H. M. Antia & S. Basu (1994).
The input physics used by the six teams also differed. Treatments of atmospheric boundary conditions included the Eddington (A. S. Eddington, 1926) and Krishna-Swamy (K. S. Krishna Swamy, 1966) $T$-$\tau$ relations, and pre-computed photosphere tables (P. H. Hauschildt et al., 1999a, b; F. Castelli & R. L. Kurucz, 2003; R. Trampedach et al., 2014). Choices of nuclear reaction rates varied from R. H. Cyburt et al. (2010), C. Angulo et al. (1999), A. Formicola et al. (2004), J. W. Hammer et al. (2005), and E. G. Adelberger et al. (1998). The mixing length formulations included L. Henyey et al. (1965), J. P. Cox & R. T. Giuli (1968) and E. Böhm-Vitense (1958), with some teams using solar-calibrated mixing length parameters, others treating it as a free variable, and still others using calibrated prescriptions for varying it. Two main metal mixtures were used: those of M. Asplund et al. (2009) and of N. Grevesse & A. J. Sauval (1998). Mass, metallicity, and initial helium abundance were all treated as free parameters by all teams.
Different teams employed different correction procedures for near-surface modeling errors. Several of the teams used empirical correction formulae (H. Kjeldsen et al., 2008; W. H. Ball & L. Gizon, 2014), where an offset is applied to the mode frequencies to account for the effects of the surface term, before the corrected mode frequencies are used to construct a conventional likelihood function. These empirical formulae admit additional free parameters, which may either be fitted against the data, or potentially be externally calibrated against 3D models (T. Sonoi et al., 2015) or from an empirical ensemble (Y. Li et al., 2023a). Other teams employed the nonparametric $\epsilon$-matching technique (I. W. Roxburgh, 2016), where combinations of mode frequencies are computed from both the model and observed set, chosen so that the likelihood function evaluated from these combinations is insensitive to the near-surface structure of the star (J. M. J. Ong et al., 2021). Table 6 lists the detailed configurations as well as the derived stellar parameters by each team.
@@
III.2 Rotational Modeling
To model the star based on its rotation period using gyrochronology, we require rotation period predictions from stellar models. For this purpose, we used Team 1’s stellar models in combination with the rotevol code (J. L. van Saders & M. H. Pinsonneault, 2013; G. Somers et al., 2017). Angular momentum loss was modeled using a calibrated prescription, based on open cluster data and asteroseismic field stars, following the exact configuration described in Y. Li et al. (2025).
We briefly summarize the calibration procedure used in the angular momentum loss modeling. Stellar models were initialized with a rotation period of $P_{\rm disk}=4$ days and a disk-locking timescale of $\tau_{\rm disk}=10$ Myr, following the setup of F. Chiti et al. (2024). We note that these initial conditions does not influence the rotation–age relation once convergence has been achieved. Angular momentum loss was modeled using the prescription from J. L. van Saders et al. (2016), which relates the loss rate to the stellar rotation rate and convective overturn timescale, scaled by a braking efficiency parameter $f_{K}$. When the Rossby number exceeds a critical threshold ${{\rm Ro}}_{\rm crit}$, the star enters a weakened magnetic braking regime, and angular momentum loss is no longer applied. This is less relevant in our case because the star has not entered into that phase. Stellar rotation was modeled using a two-zone framework in which the radiative core and convective envelope rotate as solid bodies but exchange angular momentum over a coupling timescale $\tau_{\rm ce}$. This timescale is assumed to scale with stellar mass as $\tau_{\rm ce}/\tau_{\rm ce,\odot}=(M/M_{\odot})^{-\alpha_{\rm ce}}$, where $\tau_{\rm ce,\odot}\approx 22$ Myr (F. Spada & A. C. Lanzafame, 2020). The free parameters in the model — $f_{K}$, ${{\rm Ro}}_{\rm crit}$, and $\alpha_{\rm ce}$ — were calibrated to reproduce the rotation periods of stars in a set of well-characterized open clusters ranging from the 120 Myr Pleiades to the 4 Gyr M67 (L. M. Rebull et al., 2016; S. T. Douglas et al., 2017, 2019; J. L. Curtis et al., 2019; S. Meibom et al., 2015; J. L. Curtis et al., 2020; S. A. Barnes et al., 2016; R. Dungee et al., 2022; L. Long et al., 2023), asteroseismic field stars (O. J. Hall et al., 2021; V. Silva Aguirre et al., 2015, 2017), and the Sun (J. N. Bahcall et al., 1995) at their respective ages. The $T_{\text{eff}}$ and rotation period of HR 7672A place it in a regime where gyrochronology is well calibrated.